A sample of carbon dioxide gas at a pressure of 0.668 atm and a temperature of 20.7oC, occupies a volume of 19.1 liters. If the gas is allowed to expand at constant temperature to a volume of 24.7 liters, the pressure of the gas sample will be ? atm.
1
Expert's answer
2015-10-29T07:41:55-0400
According to Boyle-Mariott 's law, at a constant temperature work of pressure of gas upon its volume constantly: P1*V1 = P2*V2.
Then pressure of the carbon dioxide occupying the volume of 24.7 l, equals:
0.668*19.1/24.7 = 0.517 atm